can you explain this to me?
Mulitply
(5the square root [2] + 3) (the square root [2] -2 square root [3])
thank you
julie

Answer by checkley71(8403)

I AM ASSUMING THAT THESE 2 TERMS ARE BEING MULTIPLIED.
---------------------------------------------------------
(5SQRT2+3)
(SQRT2-2SQRT3) NOW WE MULTIPLY THESE TERMS (STANDARD MULTIPLICATION FORM)
-------------------------
5SQRT2*SQRT2+3*SQRT2+5SQRT2*-2SQRT3+3*-2SQRT3 NOW WE COMBINE THESE TERMS.
5(SQRT2*SQRT2)=5*2=10
3SQRT2
5(SQRT2)-2(SQRT3)=(5*-2)SQRT(2*3)=-10SQRT6
3(-2SQRT3)=-6SQRT3
NOW WE COMBINE ALL THESE TERMS:
10+3SQRT2-10SQRT6-6SQRT3 ANSWER.
--------------------------------------------
BECAUSE THESE SQRT VALUES ARE DIFFERENT(2,6 & 3) WE CANNOT COMBINE THEM.